### v2

# Artem Lopatin Artem Lopatin Universidade Estadual de Campinas (UNICAMP), 651 Sergio Buarque de Holanda, 13083-859 Campinas, SP, Brazil

Abstract.
Working over the split octonions over an algebraically closed field, we solve all polynomial equations in which all the coefficients but the constant term are scalar. As a consequence, we calculate the $n^{\rm-th}$ roots of an octonion.
Keywords: polynomial equations, octonions, positive characteristic.
2020 MSC: 17A75, 17D05, 20G41.
The first author was supported by FAPESP 2023/17918-2. The second author was supported by a state assignment for the Sobolev Institute of Mathematics, Siberian Branch, Russian Academy of Sciences, FWNF-2022-0003.
1. Introduction
Assume that $\mathbb{F}$ is a field of an arbitrary characteristic $p=\mathop{\rm char}\mathbb{F}\geq 0$. All vector spaces and algebras are over $\mathbb{F}$.
The problem of solving polynomial equations was historically considered as one of key problems in mathematics, which influenced the creation of Algebraic Geometry and other branches of mathematics. Polynomial equations were considered not only over fields, but also over matrix algebras, algebras of quaternions, octonions, etc. Rodríguez-Ordóñez [30] proved that every polynomial equation over the algebra $\mathbf{A}_{\mathbb{R}}$ of Cayley numbers (i.e., the division algebra of real octonions) of positive degree with the only term of the highest degree has a solution. An explicit algorithm for a solution of the quadratic equations $x^{2}+bx+c=0$ over $\mathbf{A}_{\mathbb{R}}$ was obtained by Wag, Zhang and Zhang [32] together with criterions whether this equation has one, two or infinitely many solutions.
In general, an octonion algebra $\mathbf{C}$ (or a Cayley algebra) over the field $\mathbb{F}$ is a non-associative alternative unital algebra of dimension $8$, endowed with non-singular quadratic multiplicative form $n:\mathbf{C}\to\mathbb{F}$, which is called the norm. The norm $n$ is called isotropic if $n(a)=0$ for some non-zero $a\in\mathbf{C}$, otherwise the norm $n$ is anisotropic. In case $n$ is anisotropic, the octonion algebra $\mathbf{C}$ is a division algebra. In case $n$ is isotropic, there exists a unique octonion algebra $\mathbf{O}_{\mathbb{F}}$ over $\mathbb{F}$ with isotropic norm (see Theorem 1.8.1 of [31]). This algebra is called the split octonion algebra. Note that, if $\mathbb{F}$ is algebraically closed, then any octonion algebra is isomorphic to the split octonion algebra $\mathbf{O}_{\mathbb{F}}$ (for example, see Lemma 2.2 from [28]). Since Artin’s theorem claims that in an alternative algebra every subalgebra generated by two elements is associative, then any octonion algebra is power-associative, i.e., the subalgebra generated by a single element is associative. Therefore, given $a\in\mathbf{C}$, we can write down $a^{n}$ without specifying the brackets in the product.
Flaut and Shpakivskyi [17] considered the equation $x^{n}=a$ over real octonion division algebras. For an octonion division algebra $\mathbf{C}$ over an arbitrary field $\mathbb{F}$, Chapman [7] presented a complete method for finding the solutions of the polynomial equation $a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}=0$ over $\mathbf{C}$. Moreover, Chapman and Vishkautsan [9], working over a division algebra $\mathbf{C}$, determined the solutions of the polynomial equation $(a_{n}c)x^{n}+(a_{n-1}c)x^{n-1}+\cdots+(a_{1}c)x+(a_{0}c)=0$ and discussed the solutions of the polynomial equation $(ca_{n})x^{n}+(ca_{n-1})x^{n-1}+\cdots+(ca_{1})x+(ca_{0})=0$. Chapman and Levin [8] described a method for finding so-called “alternating roots” of polynomials over an arbitrary division Cayley-Dickson algebra. Chapman and Vishkautsan [10] examined the conditions under which, for a root $a$ of a polynomial $f(x)$ over a general Cayley–Dickson algebra, there exists a factorization $f(x)=g(x)(x-a)$ for some polynomial $g(x)$. The case of polynomial equations over an arbitrary algebra has recently been considered by Illmer and Netzer in [22], where some conditions were determined that imply the existence of a common solution of $n$ polynomial equations in $n$ variables, with an application to polynomial equations over $\mathbf{A}_{\mathbb{R}}$.
Assume that $\mathcal{A}$ is a unital algebra over $\mathbb{F}$. Consider a general polynomial equation $f(a_{1},\ldots,a_{m},x)=0$ over $\mathcal{A}$, i.e., $f(a_{1},\ldots,a_{m},x)$ is an element of the absolutely free unital algebra $\mathbb{F}\langle a_{1},\ldots,a_{m},x\rangle$ with free generators $a_{1},\ldots,a_{m},x$. Here $x\in\mathcal{A}$ is a variable and $a_{1},\ldots,a_{m}\in\mathcal{A}$ are coefficients. We aim to describe a method for calculating $x$ when $a_{1},\ldots,a_{m}$ are given. Note that we can also consider $f(a_{1},\ldots,a_{m},x)$ as an element of the relatively free algebra $\mathbb{F}_{\mathcal{A}}\langle a_{1},\ldots,a_{m},x\rangle:=\mathbb{F}\langle a_{1},\ldots,a_{m},x\rangle/{{\rm Id}({\mathcal{A}})}$ for $\mathcal{A}$, where ${{\rm Id}({\mathcal{A}})}$ stands for the T-ideal of all polynomial identities for $\mathcal{A}$. By knowing an $\mathbb{F}$-basis for $\mathbb{F}_{\mathcal{A}}\langle a_{1},\ldots,a_{m},x\rangle$, we can obtain a canonical form for $f(x)$. Therefore, a problem of solving a polynomial equation over an algebra $\mathcal{A}$ is tightly connected with the problem of explicit description of polynomial identities for $\mathcal{A}$. Although the theory of algebras with polynomial identities is a well-developed area of algebra with many deep results, there are still few results offering an explicit description of generators of T-ideals of polynomial identities for particular finite-dimensional algebras (see [14, 11, 12, 13, 16] for recent results). Due to the difficulty of solving a general equation $f(a_{1},\ldots,a_{m},x)=0$ over $\mathcal{A}$, it may be interesting to consider this equation over some vector subspace $\mathcal{V}\subset\mathcal{A}$ generating the algebra $\mathcal{A}$, i.e., to assume that $a_{1},\ldots,a_{m},x\in\mathcal{V}$. In this case, instead of polynomial identities, we should use so-called weak polynomial identities for the pair $(\mathcal{A},\mathcal{V})$ (see survey [15] and papers [23, 25, 26, 27] for recent results on weak polynomial identities).
The split-octonions have numerous applications to physics. As an example, the Dirac equation, which describes the motion of a free spin $1/2$ particle, such as an electron or a proton, can be represented by the split-octonions (see [18, 19, 20]). There exist applications of split-octonions to electromagnetic theory (see [4, 6, 5]), geometrodynamics (see [3]), unified quantum theories (see [24, 1, 2]), special relativity (see [21]).
In a recent paper, Lopatin and Zubkov [28] considered the linear equations $ax=c$, $(ax)b=c$, $a(bx)=c$ over the split octonion algebra $\mathbf{O}$ in case $\mathbb{F}$ is algebraically closed. Note that, over a division octonion algebra, these equations can easily be solved, and for non-zero $a,b$ there is a unique solution. However, in the case of an algebraically closed field, the situation is drastically different. Specifically, the set $X$ of all solutions of one of the above linear equations with non-zero $a,b$ is either empty, or contains the only element, or the affine variety $X$ has dimension $r$, where
  1. $\bullet$
    $r=4$ in case we consider the equation $ax=c$;
  2. $\bullet$
    $r\in\{4,5,7\}$ in case we consider the equation $(ax)b=c$;
  3. $\bullet$
    $r\in\{4,6,8\}$ in case we consider the equation $a(bx)=c$.
In Sections 2 and 3 we assume that the field $\mathbb{F}$ is algebraically closed. In Sections 2 we explicitly define the octonion algebra $\mathbf{O}$ and its group of automorphisms $\mathop{\rm Aut}(\mathbf{O})={\rm G}_{2}$. In Section 3 we solve every equation
$\alpha_{n}x^{n}+\alpha_{n-1}x^{n-1}+\cdots+\alpha_{1}x=c$ (1.1)
with scalar $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{F}$ and possibly non-scalar constant term $c\in\mathbf{O}$ with respect to the variable $x\in\mathbf{O}$ (see Theorem 3.2). The solution of equation (1.1) is obtained modulo solution of polynomial equations over $\mathbb{F}$. In Corollary 3.3 we explicitly describe the number of solutions of equation (1.1). In Corollary 3.4 we apply the obtained general result to the $n^{\rm-th}$ roots of $c\in\mathbf{O}$, i.e., to the solutions of the equation $x^{n}=c$.
2. Octonions
In this section we assume that the field $\mathbb{F}$ is algebraically closed.
2.1. Split-octonions
The split octonion algebra $\mathbf{O}=\mathbf{O}(\mathbb{F})$, also known as the split Cayley algebra, is the vector space of all matrices
$a=\left(\begin{array}[]{cc}\alpha&\mathbf{u}\\ \mathbf{v}&\beta\\ \end{array}\right)\text{ with }\alpha,\beta\in\mathbb{F}\text{ and }\mathbf{u},\mathbf{v}\in\mathbb{F}^{3},$
together with the multiplication given by the following formula:
$aa^{\prime}=\left(\begin{array}[]{cc}\alpha\alpha^{\prime}+\mathbf{u}\cdot\mathbf{v}^{\prime}&\alpha\mathbf{u}^{\prime}+\beta^{\prime}\mathbf{u}-\mathbf{v}\times\mathbf{v}^{\prime}\\ \alpha^{\prime}\mathbf{v}+\beta\mathbf{v}^{\prime}+\mathbf{u}\times\mathbf{u}^{\prime}&\beta\beta^{\prime}+\mathbf{v}\cdot\mathbf{u}^{\prime}\\ \end{array}\right),\text{ where }a^{\prime}=\left(\begin{array}[]{cc}\alpha^{\prime}&\mathbf{u}^{\prime}\\ \mathbf{v}^{\prime}&\beta^{\prime}\\ \end{array}\right),$
[…]
$e_{1}=\left(\begin{array}[]{cc}1&\mathbf{0}\\ \mathbf{0}&0\\ \end{array}\right),\;e_{2}=\left(\begin{array}[]{cc}0&\mathbf{0}\\ \mathbf{0}&1\\ \end{array}\right),\;\mathbf{u}_{i}=\left(\begin{array}[]{cc}0&\mathbf{c}_{i}\\ \mathbf{0}&0\\ \end{array}\right),\;\mathbf{v}_{i}=\left(\begin{array}[]{cc}0&\mathbf{0}\\ \mathbf{c}_{i}&0\\ \end{array}\right)$
for $i=1,2,3$. The unity of $\mathbf{O}$ is denoted by $1_{\mathbf{O}}=e_{1}+e_{2}$. We identify octonions
$\alpha 1_{\mathbf{O}},\;\left(\begin{array}[]{cc}0&\mathbf{u}\\ \mathbf{0}&0\\ \end{array}\right),\;\left(\begin{array}[]{cc}0&\mathbf{0}\\ \mathbf{v}&0\\ \end{array}\right)$
with $\alpha\in\mathbb{F}$, $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$, respectively. Note that $\mathbf{u}_{i}\mathbf{u}_{j}=(-1)^{\epsilon_{ij}}\mathbf{v}_{k}$ and $\mathbf{v}_{i}\mathbf{v}_{j}=(-1)^{\epsilon_{ji}}\mathbf{u}_{k}$, where $\{i,j,k\}=\{1,2,3\}$ and $\epsilon_{ij}$ is the parity of the permutation $\left(\begin{array}[]{ccc}\!1\!&\!2\!&\!3\!\\ \!k\!&\!i\!&\!j\!\\ \end{array}\right)$.
The algebra $\mathbf{O}$ is endowed with the linear involution
$\overline{a}=\left(\begin{array}[]{cc}\beta&-\mathbf{u}\\ -\mathbf{v}&\alpha\\ \end{array}\right),$
which satisfies the equality $\overline{aa^{\prime}}=\overline{a^{\prime}}\overline{a}$, a norm $n(a)=a\overline{a}=\alpha\beta-\mathbf{u}\cdot\mathbf{v}$, and a non-degenerate symmetric bilinear form $q(a,a^{\prime})=n(a+a^{\prime})-n(a)-n(a^{\prime})=\alpha\beta^{\prime}+\alpha^{\prime}\beta-\mathbf{u}\cdot\mathbf{v}^{\prime}-\mathbf{u}^{\prime}\cdot\mathbf{v}$. Define the linear function trace by $\mathop{\rm tr}(a)=a+\overline{a}=\alpha+\beta$. The subspace of traceless octonions is denoted by $\mathbf{O}_{0}=\{a\in\mathbf{O}\,|\,\mathop{\rm tr}(a)=0\}$ and the affine variety of octonions with zero norm is denoted by $\mathbf{O}_{\#}=\{a\in\mathbf{O}\,|\,n(a)=0\}$. Notice that
$\mathop{\rm tr}(aa^{\prime})=\mathop{\rm tr}(a^{\prime}a)\text{ and }n(aa^{\prime})=n(a)n(a^{\prime}).$ (2.1)
The next quadratic equation holds:
$a^{2}-\mathop{\rm tr}(a)a+n(a)=0.$ (2.2)
The algebra $\mathbf{O}$ is a simple alternative algebra, i.e., the following identities hold for $a,b\in\mathbf{O}$:
$a(ab)=(aa)b,\;\;(ba)a=b(aa).$ (2.3)
Moreover,
$\overline{a}(ab)=n(a)b,\;\;(ba)\overline{a}=n(a)b.$ (2.4)
The following remark is well-known and can easily be proven.
Remark 2.1.
Given an $a\in\mathbf{O}$, one of the following cases holds:
  1. $\bullet$
    if $n(a)\neq 0$, then there exist unique $b,c\in\mathbf{O}$ such that $ba=1_{\mathbf{O}}$ and $ac=1_{\mathbf{O}}$; moreover, in this case we have $b=c=\overline{a}/n(a)$ and we denote $a^{-1}:=b=c$.
  2. $\bullet$
    if $n(a)=0$, then $a$ does not have a left inverse as well as a right inverse.
Equalities (2.4) imply that for $a\not\in\mathbf{O}_{\#}$ we have
$a^{-1}(ab)=b,\;\;(ba)a^{-1}=b.$ (2.5)
2.2. The group ${\rm G}_{2}$
The group $\mathop{\rm Aut}(\mathbf{O})$ of all automorphisms of the algebra $\mathbf{O}$ is the exceptional simple group ${\rm G}_{2}={\rm G}_{2}(\mathbb{F})$. The group ${\rm G}_{2}$ contains a Zariski closed subgroup ${\rm SL}_{3}={\rm SL}_{3}(\mathbb{F})$. Namely, every $g\in{\rm SL}_{3}$ defines the following automorphism of $\mathbf{O}$:
$a\to\left(\begin{array}[]{cc}\alpha&\mathbf{u}g\\ \mathbf{v}g^{-T}&\beta\\ \end{array}\right),$
where $g^{-T}$ stands for $(g^{-1})^{T}$ and $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$ are considered as row vectors. For every $\mathbf{u},\mathbf{v}\in\mathbf{O}$ define $\delta_{1}(\mathbf{u}),\delta_{2}(\mathbf{v})$ from $\mathop{\rm Aut}(\mathbf{O})$ as follows:
$\delta_{1}(\mathbf{u})(a^{\prime})=\left(\begin{array}[]{cc}\alpha^{\prime}-\mathbf{u}\cdot\mathbf{v}^{\prime}&(\alpha^{\prime}-\beta^{\prime}-\mathbf{u}\cdot\mathbf{v}^{\prime})\mathbf{u}+\mathbf{u}^{\prime}\\ \mathbf{v}^{\prime}-\mathbf{u}^{\prime}\times\mathbf{u}&\beta^{\prime}+\mathbf{u}\cdot\mathbf{v}^{\prime}\\ \end{array}\right),$
$\delta_{2}(\mathbf{v})(a^{\prime})=\left(\begin{array}[]{cc}\alpha^{\prime}+\mathbf{u}^{\prime}\cdot\mathbf{v}&\mathbf{u}^{\prime}+\mathbf{v}^{\prime}\times\mathbf{v}\\ (-\alpha^{\prime}+\beta^{\prime}-\mathbf{u}^{\prime}\cdot\mathbf{v})\mathbf{v}+\mathbf{v}^{\prime}&\beta^{\prime}-\mathbf{u}^{\prime}\cdot\mathbf{v}\\ \end{array}\right).$
The group ${\rm G}_{2}$ is generated by ${\rm SL}_{3}$ and $\delta_{1}(t\mathbf{u}_{i}),\delta_{2}(t\mathbf{v}_{i})$ for all $t\in\mathbb{F}$ and $i=1,2,3$ (for example, see Section 3 of [33]). By straightforward calculations, it is easy to see that
$\hbar:\mathbf{O}\to\mathbf{O},\text{ defined by }a\to\left(\begin{array}[]{cc}\beta&-\mathbf{v}\\ -\mathbf{u}&\alpha\\ \end{array}\right),$ (2.6)
belongs to ${\rm G}_{2}$.
The action of ${\rm G}_{2}$ on $\mathbf{O}$ satisfies the next properties:
$\overline{ga}=g\overline{a},\;\mathop{\rm tr}(ga)=\mathop{\rm tr}(a),\;n(ga)=n(a),\;q(ga,ga^{\prime})=q(a,a^{\prime}).$
Thus, ${\rm G}_{2}$ acts also on $\mathbf{O}_{0}$ and $\mathbf{O}_{\#}$. The group ${\rm G}_{2}$ acts diagonally on the vector space $\mathbf{O}^{n}=\mathbf{O}\oplus\cdots\oplus\mathbf{O}$ ($n$ times) by $g(a_{1},\ldots,a_{n})=(ga_{1},\ldots,ga_{n})$ for all $g\in{\rm G}_{2}$ and $a_{1},\ldots,a_{n}\in\mathbf{O}$.
We fix a binary relation $<$ on the field $\mathbb{F}$ such that for each pair $\alpha,\beta\in\mathbb{F}$ with $\alpha\neq\beta$ exactly one of $\alpha<\beta$ or $\beta<\alpha$ holds. Note that, we do not assume that $<$ is transitive, and we do not assume compatibility with the field operations.
Proposition 2.2 (Part 1 of Proposition 3.3 from [29]).
The following set is a minimal set of representatives of ${\rm G}_{2}$-orbits on $\mathbf{O}$:
  1. 1.
    $\alpha 1_{\mathbf{O}}$,
  2. 2.
    $\left(\begin{array}[]{cc}\alpha_{1}&\mathbf{0}\\ \mathbf{0}&\alpha_{2}\\ \end{array}\right)$ with $\alpha_{1}<\alpha_{2}$,
  3. 3.
    $\left(\begin{array}[]{cc}\alpha&(1,0,0)\\ \mathbf{0}&\alpha\\ \end{array}\right)$,
where $\alpha,\alpha_{1},\alpha_{2}\in\mathbb{F}$. In other words, $\mathbf{O}$ is the disjoint union of the following ${\rm G}_{2}$-orbits:
$\alpha 1_{\mathbf{O}},\;\;O_{2}(\alpha_{1},\alpha_{2}):={\rm G}_{2}(\alpha_{1}e_{1}+\alpha_{2}e_{2}),\;\;O_{3}(\alpha):={\rm G}_{2}(\alpha 1_{\mathbf{O}}+\mathbf{u}_{1}),$
where $\alpha_{1}<\alpha_{2}$, $\alpha,\alpha_{1},\alpha_{2}\in\mathbb{F}$.
Definition 2.3.
  1. $\bullet$
    The elements from Proposition 2.2 are called canonical octonions.
  2. $\bullet$
    Given $a\in\mathbf{O}$, the diagonal elements of the canonical octonion from ${\rm G}_{2}a$ are called eigenvalues $(\alpha_{1},\alpha_{2})\in\mathbb{F}^{2}$ for $a$, where $\alpha_{1}\leq\alpha_{2}$. Note that both eigenvalues are solutions of the equation $\alpha^{2}-\mathop{\rm tr}(a)\alpha+n(a)=0$.
The following remark is a consequence of Part 2 of Proposition 3.3 from [29].
Remark 2.4.
Assume $\alpha,\alpha_{i},\beta,\beta_{i},\gamma_{i}\in\mathbb{F}$ for $i=1,2$.
1. Two octonions $\alpha_{1}1_{\mathbf{O}}+\beta_{1}\mathbf{u}_{1}$ and $\alpha_{2}1_{\mathbf{O}}+\beta_{2}\mathbf{u}_{2}$ belong to the same ${\rm G}_{2}$-orbit on $\mathbf{O}$ if and only if
  1. $\bullet$
    $\alpha_{1}=\alpha_{2}$,
  2. $\bullet$
    either $\beta_{1}=\beta_{2}=0$, or $\beta_{1}$, $\beta_{2}$ are non-zero.
2. If $\gamma_{1}\neq\gamma_{2}$, then $\alpha 1_{\mathbf{O}}+\beta\mathbf{u}_{1}\not\in O_{2}(\gamma_{1},\gamma_{2})$.
3. Polynomial equations
In this section we assume that the field $\mathbb{F}$ is algebraically closed. Given a commutative associative polynomial $f(\xi)=\alpha_{n}\xi^{n}+\cdots+\alpha_{1}\xi+\alpha_{0}\in\mathbb{F}[\xi]$, where $\alpha_{0},\ldots,\alpha_{n}\in\mathbb{F}$, we write $f^{\prime}(\xi)$ for its derivative. For each $x\in\mathbf{O}$ we naturally define the substitution $f(x)=\alpha_{n}x^{n}+\cdots+\alpha_{1}x+\alpha_{0}1_{\mathbf{O}}\in\mathbf{O}$.
Assume that $f(\xi)\in\mathbb{F}[\xi]$ is a non-zero polynomial without constant term and $\gamma\in\mathbb{F}$. The multiplicity of a root $\xi_{1}\in\mathbb{F}$ for the equation $f(\xi)=\gamma$, where $\xi\in\mathbb{F}$ is a variable, is $k>0$ such that $f(\xi)-\gamma=(\xi-\xi_{1})^{k}g(\xi)$ for some $g(\xi)\in\mathbb{F}[\xi]$ with $g(\xi_{1})\neq 0$. If $k=1$, then root $\xi_{1}$ is called simple. If $k\geq 2$, then the root $\xi_{1}$ is called multiple. It is trivial that
  1. $\bullet$
    $\xi_{1}$ is a simple root if and only if $f(\xi_{1})=\gamma$ and $f^{\prime}(\xi_{1})\neq 0$;
  2. $\bullet$
    $\xi_{1}$ is a multiple root if and only if $f(\xi_{1})=\gamma$ and $f^{\prime}(\xi_{1})=0$.
Lemma 3.1.
Assume that $f(\xi)\in\mathbb{F}[\xi]$ and $\alpha,\beta\in\mathbb{F}$. Then $f(\alpha 1_{\mathbf{O}}+\beta\mathbf{u}_{1})=f(\alpha)1_{\mathbf{O}}+f^{\prime}(\alpha)\beta\mathbf{u}_{1}$.
Proof.
For short, denote $a=\alpha 1_{\mathbf{O}}+\beta\mathbf{u}_{1}$.
1. Assume that $f(\xi)=\xi^{n}$ for some $n>0$. We prove by induction on $n$ that
$a^{n}=\alpha^{n}1_{\mathbf{O}}+n\alpha^{n-1}\beta\mathbf{u}_{1}.$ (3.1)
In case $n=1$ claim (3.1) is trivial.
Assume that claim (3.1) holds for some $n>1$. Then
$a^{n+1}=(\alpha^{n}1_{\mathbf{O}}+n\alpha^{n-1}\beta\mathbf{u}_{1})(\alpha 1_{\mathbf{O}}+\beta\mathbf{u}_{1})=\alpha^{n+1}1_{\mathbf{O}}+n\alpha^{n}\beta\mathbf{u}_{1}+\alpha^{n}\beta\mathbf{u}_{1}=\\ =\alpha^{n+1}1_{\mathbf{O}}+(n+1)\alpha^{n}\beta\mathbf{u}_{1}.$
Therefore, claim (3.1) holds for every $n>0$.
2. Assume that $f(\xi)=\alpha_{n}\xi^{n}+\cdots+\alpha_{1}\xi+\alpha_{0}\in\mathbb{F}[\xi]$ for some $\alpha_{0},\ldots,\alpha_{n}\in\mathbb{F}$ and $n\geq 0$. Note that in case $n=0$ we have $f(a)=\alpha_{0}1_{\mathbf{O}}$ and the claim of the lemma holds.
For $n>0$ we apply part 1 to obtain that
$f(a)=\sum_{i=1}^{n}\alpha_{i}a^{i}+\alpha_{0}1_{\mathbf{O}}=\sum_{i=1}^{n}\alpha_{i}(\alpha^{i}1_{\mathbf{O}}+i\alpha^{i-1}\beta\mathbf{u}_{1})+\alpha_{0}1_{\mathbf{O}},$
and the required statement is proven. ∎
Theorem 3.2.
Assume that $f(\xi)\in\mathbb{F}[\xi]$ is a non-zero polynomial without constant term and $c\in\mathbf{O}$. Acting by ${\rm G}_{2}$ on the equation $f(x)=c$, where $x\in\mathbf{O}$ is a variable, we can assume that $c$ is a canonical octonion from Proposition 2.2. Let $X\subset\mathbf{O}$ be the set of all solutions of the equation $f(x)=c$. Then
  1. 1.
    in case $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$, we have
    $\begin{array}[]{rl}X=&\big{\{}\xi_{1}1_{\mathbf{O}}\,\big{|}\,\xi_{1}\in\mathbb{F}\text{ satisfies }f(\xi_{1})=\gamma\big{\}}\bigcup\\ &\big{\{}O_{2}(\xi_{1},\xi_{2})\,\big{|}\,\xi_{1},\xi_{2}\in\mathbb{F}\text{ satisfy }f(\xi_{1})=f(\xi_{2})=\gamma\text{ and }\xi_{1}<\xi_{2}\big{\}}\bigcup\\ &\big{\{}O_{3}(\xi_{1})\,\big{|}\,\xi_{1}\in\mathbb{F}\text{ is a multiple root for }f(\xi)=\gamma\big{\}};\\ \end{array}$
  2. 2.
    in case $c=\gamma_{1}e_{1}+\gamma_{2}e_{2}$ for some $\gamma_{1},\gamma_{2}\in\mathbb{F}$ with $\gamma_{1}<\gamma_{2}$, we have
    $X=\big{\{}\xi_{1}e_{1}+\xi_{2}e_{2}\,\big{|}\,\xi_{1},\xi_{2}\in\mathbb{F}\text{ satisfy }f(\xi_{1})=\gamma_{1}\text{ and }f(\xi_{2})=\gamma_{2}\big{\}};$
  3. 3.
    in case $c=\gamma 1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\gamma\in\mathbb{F}$, we have
    $X=\Bigg{\{}\xi_{1}1_{\mathbf{O}}+\frac{1}{f^{\prime}(\xi_{1})}\mathbf{u}_{1}\,\Bigg{|}\,\xi_{1}\in\mathbb{F}\text{ is a simple root for }f(\xi)=\gamma\Bigg{\}}.$
Proof.
Assume that there exists some $x\in X$. Consider $g\in{\rm G}_{2}$ such that $gx$ is canonical. Note that
$f(gx)=gc.$ (3.2)
1. Assume that $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$. Then we can rewrite equality (3.2) as
$f(gx)=\gamma 1_{\mathbf{O}}.$ (3.3)
One of the following possibilities holds:
  1. (a)
    $gx=\xi_{1}1_{\mathbf{O}}$ for some $\xi_{1}\in\mathbb{F}$. Equality (3.3) is equivalent to $f(\xi_{1})=\gamma$. Thus, $x=g^{-1}\xi_{1}1_{\mathbf{O}}=\xi_{1}1_{\mathbf{O}}$.
  2. (b)
    $gx=\xi_{1}e_{1}+\xi_{2}e_{2}$ for some $\xi_{1},\xi_{2}\in\mathbb{F}$ with $\xi_{1}<\xi_{2}$. Equality (3.3) is equivalent to $f(\xi_{1})=f(\xi_{2})=\gamma$. Therefore, $x\in O_{2}(\xi_{1},\xi_{2})\subset X$.
  3. (c)
    $gx=\xi_{1}1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\xi_{1}\in\mathbb{F}$. Lemma 3.1 implies that equality (3.3) is equivalent to $f(\xi_{1})=\gamma$ and $f^{\prime}(\xi_{1})=0$. Therefore, $x\in O_{3}(\xi_{1})\subset X$.
2. Assume that $c=\gamma_{1}e_{1}+\gamma_{2}e_{2}$ for some $\gamma_{1},\gamma_{2}\in\mathbf{O}$ with $\gamma_{1}<\gamma_{2}$. We have that one of the following possibilities holds:
  1. (a)
    $gx=\xi_{1}1_{\mathbf{O}}$ for some $\xi_{1}\in\mathbb{F}$. Equality (3.2) implies that
    $f(\xi_{1})1_{\mathbf{O}}\in O_{2}(\gamma_{1},\gamma_{2});$
    a contradiction to Proposition 2.2.
  2. (b)
    $gx=\xi_{1}e_{1}+\xi_{2}e_{2}$ for some $\xi_{1},\xi_{2}\in\mathbb{F}$ with $\xi_{1}<\xi_{2}$. Equality (3.2) is equivalent to
    $f(\xi_{1})e_{1}+f(\xi_{2})e_{2}=g(\gamma_{1}e_{1}+\gamma_{2}e_{2}).$
    Therefore,
    $f(\xi_{1})e_{1}+f(\xi_{2})e_{2}\in O_{2}(\gamma_{1},\gamma_{2}).$
    Let $f(\xi_{1})<f(\xi_{2})$. Then Proposition 2.2 implies that $f(\xi_{1})=\gamma_{1}$ and $f(\xi_{2})=\gamma_{2}$. Hence, $g\in\mathop{\rm St}_{{\rm G}_{2}}(\gamma_{1}e_{1}+\gamma_{2}e_{2})$. Since $\mathop{\rm St}_{{\rm G}_{2}}(\gamma_{1}e_{1}+\gamma_{2}e_{2})={\rm SL}_{3}$ by Lemma 2.1 of [29], we have that $x=g^{-1}(\xi_{1}e_{1}+\xi_{2}e_{2})=\xi_{1}e_{1}+\xi_{2}e_{2}$.
    In case $f(\xi_{1})=f(\xi_{2})$ we obtain a contradiction with Proposition 2.2.
    Let $f(\xi_{1})>f(\xi_{2})$. The equality $\hbar(f(\xi_{1})e_{1}+f(\xi_{2})e_{2})=\hbar g(\gamma_{1}e_{1}+\gamma_{2}e_{2})$ implies $f(\xi_{2})e_{1}+f(\xi_{1})e_{2}=\hbar g(\gamma_{1}e_{1}+\gamma_{2}e_{2})$. Thus, it follows from Proposition 2.2 that $f(\xi_{1})=\gamma_{2}$ and $f(\xi_{2})=\gamma_{1}$. Hence, $\hbar g\in\mathop{\rm St}_{{\rm G}_{2}}(\gamma_{1}e_{1}+\gamma_{2}e_{2})$. Since $\mathop{\rm St}_{{\rm G}_{2}}(\gamma_{1}e_{1}+\gamma_{2}e_{2})={\rm SL}_{3}$ by Lemma 2.1 of [29], we have that $x=(\hbar g)^{-1}\hbar(\xi_{1}e_{1}+\xi_{2}e_{2})=\xi_{2}e_{1}+\xi_{1}e_{2}$.
  3. (c)
    $gx=\xi_{1}1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\xi_{1}\in\mathbb{F}$. Lemma 3.1 together with equality (3.2) implies that
    $f(\xi_{1})1_{\mathbf{O}}+f^{\prime}(\xi_{1})\mathbf{u}_{1}\in O_{2}(\gamma_{1},\gamma_{2});$
    a contradiction by part 2 of Remark 2.4.
3. Assume that $c=\gamma 1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\gamma\in\mathbb{F}$. We have that one of the following possibilities holds:
  1. (a)
    $gx=\xi_{1}1_{\mathbf{O}}$ for some $\xi_{1}\in\mathbb{F}$. Equality (3.2) implies that
    $f(\xi_{1})1_{\mathbf{O}}\in O_{3}(\gamma);$
    a contradiction to Proposition 2.2.
  2. (b)
    $gx=\xi_{1}e_{1}+\xi_{2}e_{2}$ for some $\xi_{1},\xi_{2}\in\mathbb{F}$ with $\xi_{1}<\xi_{2}$. Equality (3.2) implies that
    $f(\xi_{1})e_{1}+f(\xi_{2})e_{2}\in O_{3}(\gamma);$
    a contradiction to Proposition 2.2.
  3. (c)
    $gx=\xi_{1}1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\xi_{1}\in\mathbb{F}$. Lemma 3.1 implies that equality (3.2) is equivalent to
    $f(\xi_{1})1_{\mathbf{O}}+f^{\prime}(\xi_{1})\mathbf{u}_{1}=g(\gamma 1_{\mathbf{O}}+\mathbf{u}_{1}).$ (3.4)
    It follows from part 1 of Remark 2.4 that equality (3.4) is equivalent to $f(\xi_{1})=\gamma$, $f^{\prime}(\xi_{1})\neq 0$, and $g\mathbf{u}_{1}=f^{\prime}(\xi_{1})\mathbf{u}_{1}$. Thus,
    $x=g^{-1}(\xi_{1}1_{\mathbf{O}}+\mathbf{u}_{1})=\xi_{1}1_{\mathbf{O}}+g^{-1}\mathbf{u}_{1}=\xi_{1}1_{\mathbf{O}}+\frac{1}{f^{\prime}(\xi_{1})}\mathbf{u}_{1}.$
    On the other hand, Lemma 3.1 implies that for every $\xi_{1}\in\mathbb{F}$ with $f(\xi_{1})=\gamma$ and $f^{\prime}(\xi_{1})\neq 0$ we have
    $f\Bigg{(}\xi_{1}1_{\mathbf{O}}+\frac{1}{f^{\prime}(\xi_{1})}\mathbf{u}_{1}\Bigg{)}=\gamma 1_{\mathbf{O}}+\mathbf{u}_{1}.$
The following corollaries are immediate consequences of Theorem 3.2.
Corollary 3.3.
Assume that $f(\xi)\in\mathbb{F}[\xi]$ is a non-zero polynomial of degree $n>1$ without constant term and $c\in\mathbf{O}$. Denote by $X\subset\mathbf{O}$ the set of all solutions of the equation $f(x)=c$. Let $(\gamma_{1},\gamma_{2})\in\mathbb{F}^{2}$ be the eigenvalues of $c$. Then
  1. (a)
    $X$ is infinite if and only if $\gamma_{1}=\gamma_{2}$ and $c=\gamma_{1}1_{\mathbf{O}}$;
  2. (b)
    $X$ is empty if and only if $\gamma_{1}=\gamma_{2}$, $c\neq\gamma_{1}1_{\mathbf{O}}$, and the equation $f(\xi)=\gamma_{1}$, where $\xi\in\mathbb{F}$ is a variable, does not have any simple root;
  3. (c)
    $|X|\leq n$ in case $\gamma_{1}=\gamma_{2}$ and $c\neq\gamma_{1}1_{\mathbf{O}}$;
  4. (d)
    $1\leq|X|\leq n^{2}$ in case $\gamma_{1}\neq\gamma_{2}$.
Corollary 3.4.
Assume that $n>1$ is an integer and $c\in\mathbf{O}$. Acting by ${\rm G}_{2}$ on the equation $x^{n}=c$, where $x\in\mathbf{O}$ is a variable, we can assume that $c$ is a canonical octonion from Proposition 2.2. Let $X\subset\mathbf{O}$ be the set of all solutions of the equation $x^{n}=c$. Then
  1. 1.
    in case $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$, we have
    $\begin{array}[]{rl}X=&\big{\{}\xi_{1}1_{\mathbf{O}}\,\big{|}\,\xi_{1}\in\mathbb{F}\text{ satisfies }\xi_{1}^{n}=\gamma\big{\}}\bigcup\\ &\big{\{}O_{2}(\xi_{1},\xi_{2})\,\big{|}\,\xi_{1},\xi_{2}\in\mathbb{F}\text{ satisfy }\xi_{1}^{n}=\xi_{2}^{n}=\gamma\text{ and }\xi_{1}<\xi_{2}\big{\}}\bigcup Y,\text{ where }\\ Y=&\big{\{}O_{3}(\xi_{1})\,\big{|}\,\xi_{1}\in\mathbb{F}\text{ satisfies }\xi_{1}^{n}=\gamma\big{\}}\;\text{ if }p|n\text{ or }\gamma=0;\quad\text{and }Y=\emptyset,\;\text{ otherwise};\\ \end{array}$
  2. 2.
    in case $c=\gamma_{1}e_{1}+\gamma_{2}e_{2}$ for some $\gamma_{1},\gamma_{2}\in\mathbb{F}$ with $\gamma_{1}<\gamma_{2}$, we have
    $X=\big{\{}\xi_{1}e_{1}+\xi_{2}e_{2}\,\big{|}\,\xi_{1},\xi_{2}\in\mathbb{F}\text{ satisfy }\xi_{1}^{n}=\gamma_{1}\text{ and }\xi_{2}^{n}=\gamma_{2}\big{\}};$
  3. 3.
    in case $c=\gamma 1_{\mathbf{O}}+\mathbf{u}_{1}$ for some $\gamma\in\mathbb{F}$, we have
    $X=\Bigg{\{}\xi_{1}1_{\mathbf{O}}+\frac{\xi_{1}}{n\gamma}\mathbf{u}_{1}\,\Bigg{|}\,\xi_{1}\in\mathbb{F}\text{ satisfies }\xi_{1}^{n}=\gamma\Bigg{\}},\quad\text{ if }p\not{|}n\text{ and }\gamma\neq 0;$
    $X=\emptyset,\quad\text{ if }p|n\text{ or }\gamma=0.\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\;$
Acknowledgement
We sincerely thank the four anonymous referees for their valuable and helpful comments.